OpenStudy (anonymous):

Need help solving this please! (-6/5)x-(5/6)=(2/3)x+(3/2)

OpenStudy (anonymous):

(-6/5)x-(5/6)=(2/3)x+(3/2)<->x=-5/4

OpenStudy (anonymous):

how did you do that?can you please show me the steps?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

how did you get the 7/3 in the 3rd step?